# Probabilistic Super-Resolution for Urban Micrometeorology via a Schrödinger Bridge

### v1

@@ 1 Introduction
Deep learning-based super-resolution (SR) has been applied to accelerate numerical weather prediction [e.g., 28, 47, 27]. In this approach, trained neural networks enhance the resolution of predictions, enabling rapid high-resolution (HR) inference without costly HR numerical integration. Such acceleration has been demonstrated not only for global and mesoscale problems [e.g., 47, 27] but also for microscale problems in urban areas [28, 50, 42, 51].
@@
Recently, denoising diffusion probabilistic models [simply, diffusion models (DMs); 15] have been actively applied to SR in meteorology [20, 13, 25, 36, 43]. In DM-based SR, HR samples are generated by repeatedly transforming noise. This sequence of transformations is described by diffusion processes, mathematically formulated as stochastic differential equations (SDEs) with time tendencies parameterized by neural networks [40]. Iterative transformations with SDEs yield accurate SR inference [e.g., 34] and allow uncertainty quantification, since the generated samples follow the HR data distribution [e.g., 25].
@@
SR aims to convert low-resolution (LR) data into HR data, suggesting that DM-based SR becomes more efficient when it starts from LR inputs rather than from noise. The Schrödinger bridge (SB) provides a general framework, including DMs, for transformations between arbitrary data distributions [24]. This problem reduces to estimating SDEs; once an SDE is learned by a neural network, its integration provides the desired transformation [e.g., 8, 6]. In the context of SR, the estimated SDE directly transforms LR data into HR data, leading to greater efficiency than DMs [22].
@@
In meteorology, applications of DMs are rapidly increasing [20, 13, 25, 36, 43], whereas SB-based neural networks remain largely unexplored. In computer vision, where SB methods are advancing, inference accuracy has been widely studied [e.g., 22, 48], but inference uncertainty has received little attention. Since SB methods also reconstruct data distributions via SDEs, they should allow uncertainty quantification as in DMs. Such evaluation is essential in meteorological problems, which inherently involve uncertainty.
@@
This study applies an SB-based SR model [7] to 2-m temperature in an actual urban area. We show that the SB model achieves more efficient probabilistic SR (i.e., ensemble SR inference) than a DM and yields improved ensemble statistics.
2 Probabilistic super‑resolution (SR)
@@
We infer HR data $\bm{x}_{\mathrm{HR}}\in\mathbb{R}^{n}$ from LR data $\bm{x}_{\mathrm{LR}}\in\mathbb{R}^{n}$ and auxiliary data $\bm{\xi}\in\mathbb{R}^{m\times n}$. All data are interpolated to $n$ grid points at HR using an interpolation method, and $\bm{\xi}$ consists of $m$ variables, such as topographic height. We consider supervised learning, where $\bm{x}_{\mathrm{HR}}$ serves as the ground truth. HR outputs from neural networks are referred to as HR samples. Since SR is an inverse problem [e.g., 30], the solution is not unique, and multiple HR samples are plausible for a given $\bm{x}_{\mathrm{LR}}$. This uncertainty is represented by the conditional distribution $p_{\mathrm{HR}}(\bm{x}\mid\bm{x}_{\mathrm{LR}},\bm{\xi})$ [e.g., 20]. The SR problem is thus formulated as a probabilistic generative task, with the objective of approximating $p_{\mathrm{HR}}(\bm{x}\mid\bm{x}_{\mathrm{LR}},\bm{\xi})$ using neural networks.
@@
We employ an SB-based SR model, simply the Schrödinger-bridge model (SM; see Appendix A for details). This model, originally proposed for forecasting tasks [7], can be adapted to SR. Specifically, we obtain an SDE solution to a particular SB problem that transforms the point mass $\delta(\bm{x}-\bm{x}_{\mathrm{LR}})$ into $p_{\mathrm{HR}}(\bm{x}\mid\bm{x}_{\mathrm{LR}},\bm{\xi})$ (Fig. 1). This SDE is learned by a neural network, and its numerical integration generates HR samples directly from LR inputs.
@@
Figure 1: Schematic of data transformation via SDEs. (Top) diffusion model [DM; 15]; (Bottom) Schrödinger-bridge model [SM; 7].
@@
For comparison, we use a DM as the baseline [15]. This DM transforms a standard Gaussian ${\cal N}(\bm{0},I_{n})$ into $p_{\mathrm{HR}}(\bm{x}\mid\bm{x}_{\mathrm{LR}},\bm{\xi})$ (Appendix B). Specifically, the DM employs a forward SDE that gradually perturbs the HR data with noise (Fig. 1), and a neural network learns the corresponding reverse dynamics—strictly speaking, the score function. Integrating the learned reverse SDE generates HR samples from noise (Fig. 1). In this formulation, $\bm{x}_{\mathrm{LR}}$ is input to the neural network together with $\bm{\xi}$ as auxiliary information. Note that the SM requires only a single directional SDE [7], whereas some Schrödinger-bridge formulations require both forward and reverse SDEs [e.g., 6].
@@
SMs are considered more efficient than DMs for two reasons. First, SMs generate HR samples from $\bm{x}_{\mathrm{LR}}$, whereas DMs generate them from Gaussian noise. Compared with noise lacking spatial structure, $\bm{x}_{\mathrm{LR}}$ is expected to have spatial structure similar to that of HR samples. Thus, the transformation with the SM can be performed with fewer steps [e.g., 22]. Hereafter, we refer to this step as a diffusion-time step to distinguish it from physical time. Second, DMs do not represent exact SB solutions and, in principle, require many diffusion-time steps [8]. Specifically, both the endpoint of the forward process and the starting point of the reverse process must be effectively Gaussian [16], which demands strong relaxation to Gaussianity—well resolved only with many steps [8]. In contrast, SMs have no such constraint and are expected to attain high accuracy with fewer steps.
@@ 3 Methods
We super-resolved 2-m temperature from 20-m to 5-m resolution (hereafter LR and HR, respectively). These data were obtained from reproduction experiments of extremely hot days during 2013–2020 [51], conducted with a physics-based micrometeorological model, the Multi-Scale Simulator for the Geoenvironment [MSSG; 29, 41, 35, 26]. The LR and HR results were computed in separate simulations with common initial and boundary conditions. Unlike creating LR data by averaging HR results, this setting makes SR more difficult, since the temporal evolution within the computational domain is simulated separately for LR and HR [46, 47, 51].
@@ 3.1 Data
The dataset covers a 1.6-km square area centered on Tokyo Station (Fig. 2) and consists of MSSG outputs averaged over 1-min intervals. For details, see [51]. The LR and HR grids are $80\times 80$ at 20-m resolution and $320\times 320$ at 5-m resolution, respectively. Inputs to the DM and SM consist of LR temperature at 2 m height, together with LR temperature and LR three-dimensional velocity in the lowest seven vertical levels. Additionally, HR building height and HR land‑use index are used as static inputs. All inputs are stacked along the channel dimension and upsampled to $320\times 320$ using nearest‑neighbor interpolation. We denote the LR 2-m temperature by $\bm{x}_{\mathrm{LR}}$, and all other inputs collectively by $\bm{\xi}$. The output variable is the HR 2-m temperature. Data from 2013–2018 (2,387 sets) were used for training, data from 2019 (493 sets) for validation, and data from 2020 (540 sets) for testing. Hyperparameters were tuned with the validation data (Section 3.3). All results in Section 4 are based on the test data.
@@ 3.2 Neural networks
For the DM, we adopted the Palette model [33], a U-Net based DM designed for SR tasks. To improve accuracy, we applied residual learning [18, 25]: the DM is trained to generate $\bm{x}_{\mathrm{HR}}-\bm{x}_{\mathrm{LR}}$, and during testing, $\bm{x}_{\mathrm{LR}}$ is added to the generated samples (Appendix B). For the SM, we used the same U‑Net architecture to parameterize the SDE drift (Appendix A). We briefly summarize this U-Net below; for details, see our public code (Code Availability).
@@
Although various acceleration methods have been developed for DMs and SMs [e.g., 39, 17, 4, 48], we did not use them. For example, implicit acceleration methods (e.g., DDIM) have been proposed for both DMs and SMs [39, 48]. However, to highlight the differences between the DM and SM frameworks, we followed their original formulations [15, 7], using the same U-Net architecture.
@@ 3.3 Training
The DM and SM were trained with AdamW [23] using a learning rate of $1\times 10^{-4}$, a batch size of 32, and 1,000 epochs. The SM loss was the mean squared error (MSE) between the SDE drift and its U-Net approximation (Appendix A), whereas the DM loss was the MSE for denoising score matching (Appendix B).
The following hyperparameters were tuned to minimize the root mean square error (RMSE) on the validation data: number of epochs, learning rate, batch size, and noise amplitudes for both the DM and SM. These parameters were fixed across all experiments after tuning, and the results were not highly sensitive to their exact values. For the SM, the noise amplitude decreased from $2.0\times 10^{-1}$ to 0 with diffusion time ($\gamma_{t}$ in Appendix A). For the DM, the drift coefficient decreased linearly from $1\times 10^{1}$ to $1\times 10^{-3}$ ($\lambda_{t}$ in Appendix B), giving a maximum noise amplitude of $\sqrt{1\times 10^{1}}\approx 3.16$. We confirmed that the tuned values were comparable to those reported in previous studies [15, 33, 34, 7].
@@ 3.4 Evaluation metrics
SR accuracy was evaluated with two metrics: RMSE, which quantifies point-wise error, and the structural similarity index measure (SSIM) loss, which assesses pattern similarity. For both metrics, smaller values indicate results closer to the ground truth. These metrics are widely used in SR research [5, 19]. In practice, we generated one HR sample for each ground-truth datum, computed the metrics, and averaged them over the test data (540 sets). The mean values were nearly independent of the random noise used during sample generation.
Inference statistics were evaluated with 64-member ensemble inference for each ground-truth datum. Similar results were obtained with 32-member ensembles. For the SM, all members were initialized from the same LR state (2-m temperature), and diversity arose from stochasticity during SDE integration (Fig. 1). For the DM, each member was initialized with different Gaussian noise and further perturbed during integration (Fig. 1).
We report the spread–skill ratio (Spread/RMSE) and the rank histogram, both of which are standard diagnostics [49, 10]. The spread–skill ratio compares the ensemble spread (i.e., standard deviation) with the RMSE between the ensemble mean and the ground truth. Ratios close to 1 indicate appropriate uncertainty, whereas ratios less than 1 (or greater than 1) indicate underdispersion (or overdispersion). The rank histogram collects the ranks (1–65) of ground-truth values relative to 64 members. Flat histograms indicate reliable dispersion, whereas U-shaped histograms—often observed for deep learning models [e.g., 25]—indicate underdispersion; that is, a tendency for the ground truth to lie outside the ensemble range. While rank histograms are constructed from ranks at each grid point, the spread–skill ratio is computed from RMSEs and spreads after spatial averaging.
@@ 4.1 Accuracy for SR inference
Figure 3 shows an example of SR. Compared with the input LR field, the HR simulation resolves fine-scale buildings and the associated temperature patterns. Both the DM and SM reproduce these HR patterns well, with $N_{T}=50$ for the DM and $N_{T}=10$ for the SM (Figs. 3a–3c); thus, the SM attains comparable accuracy at approximately one-fifth the computational cost of the DM. A single ensemble member closely resembles the ensemble mean at small scales, and averaging introduces only mild blurring. This likely reflects that HR building geometry governs the spatial scales of temperature [52], and both the DM and SM use this static HR building field as an important auxiliary input. Comparing the distributions of absolute error and ensemble spread, we find that spreads tend to be larger in regions with larger errors (Figs. 3d and 3e). Indeed, the Pearson correlation between absolute error and spread is about 0.32 for both the DM and SM, which is significantly positive.
Figure 3: Example of SR for 2‑m temperature over a 500 m $\times$ 500 m area in Yaesu, Tokyo (see Fig. 2): (a) an ensemble member, (b) ensemble mean, (c) ground-truth datum (i.e., HR datum), (d) absolute error between the ensemble mean and ground truth, and (e) ensemble spread (i.e., standard deviation). The top row shows diffusion model (DM) results, and the bottom row shows Schrödinger‑bridge model (SM) results. In the middle column (c), the bottom row shows the input LR 2-m temperature. The DM uses $N_{T}=50$, whereas the SM uses $N_{T}=10$.
@@
Figure 4: Dependence of mean test errors for 2-m temperature on the number of diffusion-time steps $N_{T}$. For example, with $N_{T}=10$, the diffusion time $t\in[0,1]$ is divided into 10 steps, and the SDE is numerically integrated using the Euler-Maruyama method. Mean test errors were calculated by performing single-member inference for each ground truth and then averaged over all test data (540 sets). Error bars indicate the standard deviations over five experiments with different random initializations of the U-Nets.
@@
Table 1: Mean test errors for the DM and SM. Errors were computed by performing single-member inference for each ground truth and then averaged over all test data (540 sets). Standard deviations were obtained from five experiments with different random initializations of the U-Nets.

### v2

@@ 1 Introduction
Deep learning-based super-resolution (SR) has been applied to accelerate numerical weather prediction [e.g., 30, 49, 29]. In this approach, trained neural networks enhance the resolution of predictions, enabling rapid high-resolution (HR) inference without costly HR numerical integration. Such acceleration has been demonstrated not only for global and mesoscale problems [e.g., 49, 29] but also for microscale problems in urban areas [30, 52, 44, 53].
@@
Recently, denoising diffusion probabilistic models [simply, diffusion models (DMs); 17] have been actively applied to SR in meteorology [22, 15, 27, 38, 45]. In DM-based SR, HR samples are generated by repeatedly transforming noise. This sequence of transformations is described by diffusion processes, mathematically formulated as stochastic differential equations (SDEs) with time derivatives parameterized by neural networks [42]. Iterative transformations with SDEs yield accurate SR inference [e.g., 36] and allow uncertainty quantification, since the generated samples follow the HR data distribution [e.g., 27].
@@
SR aims to convert low-resolution (LR) data into HR data, suggesting that DM-based SR becomes more efficient when it starts from LR inputs rather than from noise. The Schrödinger bridge (SB) provides a generalization of DMs for transformations between arbitrary data distributions [26]. This problem reduces to estimating SDEs; once an SDE is learned by a neural network, its integration provides the desired transformation [e.g., 9, 6]. In the context of SR, the estimated SDE directly transforms LR data into HR data, leading to greater efficiency than DMs [24].
@@
In meteorology, applications of DMs are rapidly increasing [22, 15, 27, 38, 45], whereas SB-based neural networks remain largely unexplored. In computer vision, where SB methods are advancing, inference accuracy has been widely studied [e.g., 24, 50], but inference uncertainty has received little attention. Since SB methods also reconstruct data distributions via SDEs, they should allow uncertainty quantification as in DMs. Such evaluation is essential in meteorological problems, which inherently involve uncertainty.
@@
This study applies an SB-based SR model [7] to 2-m temperature in an actual urban area. We show that the SB model achieves more efficient probabilistic SR (i.e., ensemble SR inference) than a DM and yields improved ensemble statistics. For clarity, a list of abbreviations is provided in Table A1 of Appendix A.
2 Probabilistic super-resolution (SR)
@@
We infer HR data $\bm{x}_{\mathrm{HR}}\in\mathbb{R}^{n}$ from LR data $\bm{x}_{\mathrm{LR}}\in\mathbb{R}^{n}$ and auxiliary data $\bm{\xi}\in\mathbb{R}^{m\times n}$. All data are interpolated to $n$ grid points at HR using an interpolation method, and $\bm{\xi}$ consists of $m$ variables, such as topographic height. We consider supervised learning, where $\bm{x}_{\mathrm{HR}}$ serves as the ground truth. HR outputs from neural networks are referred to as HR samples. Since SR is an inverse problem [e.g., 32], the solution is not unique, and multiple HR samples are plausible for a given $\bm{x}_{\mathrm{LR}}$. This uncertainty is represented by the conditional distribution $p_{\mathrm{HR}}(\bm{x}\mid\bm{x}_{\mathrm{LR}},\bm{\xi})$ [e.g., 22]. The SR problem is thus formulated as a probabilistic generative task, with the objective of approximating $p_{\mathrm{HR}}(\bm{x}\mid\bm{x}_{\mathrm{LR}},\bm{\xi})$ using neural networks.
@@
We employ an SB-based SR model, simply the Schrödinger-bridge model (SM; see Appendix B for details). This model, originally proposed for forecasting tasks [7], can be adapted to SR. Specifically, we obtain an SDE solution to a particular SB problem that transforms the point mass $\delta(\bm{x}-\bm{x}_{\mathrm{LR}})$ into $p_{\mathrm{HR}}(\bm{x}\mid\bm{x}_{\mathrm{LR}},\bm{\xi})$ (Fig. 1). This SDE is learned by a neural network, and its numerical integration generates HR samples directly from LR inputs.
@@
Figure 1: Schematic of data transformation via SDEs. (Top) diffusion model [DM; 17]; (Bottom) Schrödinger-bridge model [SM; 7].
@@
For comparison, we use a DM as the baseline [17]. This DM transforms a standard Gaussian ${\cal N}(\bm{0},I_{n})$ into $p_{\mathrm{HR}}(\bm{x}\mid\bm{x}_{\mathrm{LR}},\bm{\xi})$ (see Appendix C). Specifically, the DM employs a forward SDE that gradually perturbs the HR data with noise (Fig. 1), and a neural network learns the corresponding reverse dynamics—strictly speaking, the score function. Integrating the learned reverse SDE generates HR samples from noise (Fig. 1). In this formulation, $\bm{x}_{\mathrm{LR}}$ is input to the neural network together with $\bm{\xi}$ as auxiliary information. Note that the SM requires only a single directional SDE [7], whereas some Schrödinger-bridge formulations require both forward and reverse SDEs [e.g., 6]. The behavior of the SM also depends on its reference SDE, and we summarize how this reference differs from that of the DM in Appendix D.
@@
SMs are considered more efficient than DMs for two reasons. First, SMs generate HR samples from $\bm{x}_{\mathrm{LR}}$, whereas DMs generate them from Gaussian noise. Compared with noise lacking spatial structure, $\bm{x}_{\mathrm{LR}}$ is expected to have spatial structure similar to that of HR samples. Thus, the transformation with the SM can be performed with fewer steps [e.g., 24]. Hereafter, we refer to this step as a diffusion-time step to distinguish it from physical time. Second, DMs do not represent exact SB solutions and, in principle, require many diffusion-time steps [9]. Specifically, both the endpoint of the forward process and the starting point of the reverse process must be effectively Gaussian [18], which demands strong relaxation to Gaussianity—well resolved only with many steps [9]. In contrast, SMs have no such constraint and are expected to attain high accuracy with fewer steps. Further detailed comparisons between DMs and SMs are provided in Appendix D.
@@ 3 Methods
We super-resolved 2-m temperature from 20-m to 5-m resolution (hereafter LR and HR, respectively). These data were obtained from reproduction experiments of extremely hot days during 2013–2020 [53], conducted with a physics-based micrometeorological model, the Multi-Scale Simulator for the Geoenvironment [MSSG; 31, 43, 37, 28]. The LR and HR results were computed in separate simulations with common initial and boundary conditions. Unlike creating LR data by averaging HR results, this setting makes SR more difficult, since the temporal evolution within the computational domain is simulated separately for LR and HR [48, 49, 53].
@@ 3.1 Data
The datasets cover a 1.6-km square area centered on Tokyo Station (Fig. 2) and consist of MSSG outputs, i.e., 1-min averaged temperature and velocity fields sampled at 1-min intervals. We define one set (i.e., one dataset) as a pair of LR and HR fields at a single time step. A temporal resolution of 1 min is sufficient to characterize mean flows in street canyons [e.g., 8]. For details, see [53]. The LR and HR grids are $80\times 80$ at 20-m resolution and $320\times 320$ at 5-m resolution, respectively. Inputs to the DM and SM consist of LR temperature at 2 m height, together with LR temperature and LR three-dimensional velocity in the lowest seven vertical levels. Additionally, HR building height and HR land-use index are used as static inputs. All inputs are stacked along the channel dimension and upsampled to $320\times 320$ using nearest-neighbor interpolation. We denote the LR 2-m temperature by $\bm{x}_{\mathrm{LR}}$, and all other inputs collectively by $\bm{\xi}$. The output variable is the HR 2-m temperature. Data from 2013–2018 (2,387 sets) were used for training, data from 2019 (493 sets) for validation, and data from 2020 (540 sets) for testing. Hyperparameters were tuned with the validation data (Section 3.3). All results in Section 4 are based on the test data. We confirmed that similar results are obtained when using a random data partitioning that ignores temporal order, indicating that the results are not sensitive to the choice of data partitioning.
@@ 3.2 Neural networks
For the DM, we adopted the Palette model [35], a U-Net based DM designed for SR tasks. To improve accuracy, we applied residual learning [20, 27]: the DM is trained to generate $\bm{x}_{\mathrm{HR}}-\bm{x}_{\mathrm{LR}}$, and during testing, $\bm{x}_{\mathrm{LR}}$ is added to the generated samples. For the SM, we used the same U-Net architecture to parameterize the SDE drift (see Appendix B). We briefly summarize this U-Net below; for details, see our public code (Code Availability).
@@
Although various acceleration methods have been developed for DMs and SMs [e.g., 41, 19, 4, 50], we did not use them. For example, implicit acceleration methods (e.g., DDIM) have been proposed for both DMs and SMs [41, 50]. However, to highlight the differences between the DM and SM frameworks, we followed their original formulations [17, 7], using the same U-Net architecture.
@@ 3.3 Training
The DM and SM were trained with AdamW [25] using a learning rate of $1\times 10^{-4}$, a batch size of 32, and 1,000 epochs. The SM loss was the mean squared error (MSE) between the SDE drift and its U-Net approximation, whereas the DM loss was the MSE for denoising score matching (see Appendices B and C).
The following hyperparameters were tuned to minimize the root mean square error (RMSE) on the validation data: number of epochs, learning rate, batch size, and noise amplitudes for both the DM and SM. These parameters were fixed across all experiments after tuning, and the results were not highly sensitive to their exact values. For the SM, the noise amplitude decreased from $2.0\times 10^{-1}$ to 0 with diffusion time ($\gamma_{t}$; see Appendix B). For the DM, the drift coefficient decreased linearly from $1\times 10^{1}$ to $1\times 10^{-3}$ ($\lambda_{t}$; see Appendix C), giving a maximum noise amplitude of $\sqrt{1\times 10^{1}}\approx 3.16$. We confirmed that the tuned values were comparable to those reported in previous studies [17, 35, 36, 7].
@@ 3.4 Evaluation metrics
SR accuracy was evaluated with two metrics: RMSE, which quantifies pointwise error, and the structural similarity index measure (SSIM) loss, which assesses pattern similarity. For both metrics, smaller values indicate results closer to the ground truth. These metrics are widely used in SR research [5, 21]. In practice, we generated one HR sample for each ground-truth datum, computed the metrics, and averaged them spatially and temporally over the test data (540 sets). The mean values were nearly independent of the random noise used during sample generation.
Inference statistics were evaluated with 64-member ensembles for each ground-truth datum. Similar results were obtained with 32-member ensembles. For the SM, all members were initialized from the same LR state (2-m temperature), and diversity arose from stochasticity during SDE integration (Fig. 1). For the DM, each member was initialized with different Gaussian noise and further perturbed during integration (Fig. 1).
We report the spread–skill ratio (Spread/RMSE) and the rank histogram, both of which are standard diagnostics [51, 11]. The spread–skill ratio compares the ensemble spread (i.e., standard deviation) with the RMSE between the ensemble mean and the ground truth. Ratios close to 1 indicate appropriate uncertainty, whereas ratios less than 1 (or greater than 1) indicate underdispersion (or overdispersion). The rank histogram collects the ranks (1–65) of ground-truth values relative to 64 members. Flat histograms indicate reliable dispersion, whereas U-shaped histograms—often observed for deep learning models [e.g., 27]—indicate underdispersion; that is, a tendency for the ground truth to lie outside the ensemble range. While rank histograms are constructed from ranks at each grid point, the spread–skill ratio is computed from RMSEs and spreads after spatial averaging.
@@ 4.1 Accuracy for SR inference
Figure 3 shows an example of SR. Compared with the input LR field, the HR simulation resolves fine-scale buildings and the associated temperature patterns. Both the DM and SM reproduce these HR patterns well, with $N_{T}=50$ for the DM and $N_{T}=10$ for the SM (Figs. 3a–3c); thus, the SM attains comparable accuracy at approximately one-fifth the computational cost of the DM. A single ensemble member closely resembles the ensemble mean at small scales, and averaging introduces only mild blurring. This likely reflects that HR building geometry governs the spatial scales of temperature [54], and both the DM and SM use this static HR building field as an important auxiliary input. Comparing the distributions of absolute error and ensemble spread, we find that spreads tend to be larger in regions with larger errors (Figs. 3d and 3e). Indeed, the Pearson correlation between absolute error and spread is about 0.32 for both the DM and SM, which is significantly positive. To further examine this relationship, we also evaluated RMSE and ensemble spread after excluding building areas. When focusing on ground-only regions, the correspondence between RMSE and spread becomes clearer, implying that building-induced variability can partially obscure this relationship.
Figure 3: SR for 2-m temperature over a 500 m $\times$ 500 m area in Yaesu, Tokyo (see Fig. 2) at 2020-08-15 14:46+09:00: (a) an ensemble member, (b) ensemble mean, (c) ground-truth datum (i.e., HR datum), (d) absolute error between the ensemble mean and ground truth, and (e) ensemble spread (i.e., standard deviation). The top row shows diffusion model (DM) results, and the bottom row shows Schrödinger-bridge model (SM) results. In the middle column (c), the bottom row shows the input LR 2-m temperature. The DM uses $N_{T}=50$, whereas the SM uses $N_{T}=10$. The 2-m temperature is evaluated 2 m above building surfaces, or above the ground in the absence of buildings.
@@
Figure 4: Dependence of mean test errors for 2-m temperature on the number of diffusion-time steps $N_{T}$. For example, with $N_{T}=10$, the diffusion time $t\in[0,1]$ is divided into 10 steps, and the SDE is numerically integrated using the Euler–Maruyama method. Mean test errors were calculated by performing single-member inference for each ground truth and then averaging over all test data (540 sets). Error bars indicate the standard deviations over five experiments with different random initializations of the U-Nets.
@@
Table 1: Mean test errors for the DM and SM. Errors were computed by performing single-member inference for each ground truth and then averaging spatially and temporally over all test data (540 sets). Standard deviations were obtained from five experiments with different random initializations of the U-Nets.